OpenStudy (anonymous):

A square window has a side length of 40 inches. What is the approximate length of the diagonal of the window? A. 1600 inches B. 80 inches C. 57 inches D. 40 inches

OpenStudy (anonymous):

C

OpenStudy (anonymous):

40 root 2

OpenStudy (anonymous):

You need to understand pythagora's therom to do this. Pythagora's therom states that the square of the length of the hypotenuse (the longest side) is equal to the square of the other two sides added together.This only works for right angled triangles. The forumula is: c² = a²+b² Where c = hypotenuse. So in your equation, the diagonal of the square is the hypotenuse and the sides of the square are a and b. So c² = 40² + 40² therefore c² = 1600 + 1600 therefore c = √ 3200 therefore c = 56.57 inches (rounded) So your diagonal is C, 57 inches (to the nearest inch)

OpenStudy (anonymous):

You don't necessarily need pythag, it's a special right triangle being a 45 degree 45 degree 90. with the sides being a, a, and \[a \sqrt{2}\]

OpenStudy (anonymous):

Wow, I never knew that Tommy389, thank you for sharing it! Could you please explain further about the rule (like why it's like this?) What kind of rule is it? (E.g Pythagoras....). I'm noting that down for my maths summary book - thanks so much!